Reports 1-1 of 1 Clear search Modify search
AdV-SGD (FDS commissioning)
vardaro - 1:41 Monday 25 October 2021 (53630) Print this report
IR Hand off restored

The aim of the today shift was try again to recover the stability of the IR lock of the filter cavity. 

Preparation of the system:

When I started I found a correction of a -45V in the GR_M7_Y mirror. Was a mistake left on friday at the end of the green matching shift. Then I tuned the green demodulation phase at 0.615 rad.

When I started to check the IR transmission from the filter cavity was about 1V thus I started checking the alignment.

  1. I reduced the power of the SC and BAB by rotating the EQB1 FI waveplate by 15000 steps. Then I tried to engage the SC AA Loop. It did not work thus I aligned by end the SC. I did not debugged the loop. I think the problem arised on thursday morning when I restarted the HD AA process, Fiodor two weeks ago fixed this loop but maybe some parameter was not saved in the configuration file. We will check the loop the next week. In any case the SC was quite well aligned into the OPA
  2. I started to align the cavity acting on SQB1 M21 and M22. At the end of the process I also checked the alignment sending ramps to I3 M3 and M5 mirrors. The transmission was about 10V whereas the BAB transmission was about 1.7V. I did not check the quality of the alignment searching the TEM10 mode
  3. I checked the IR PDH signal and its amplitute was 1/3 of the usual one. Thus I checked the DC value of the EQB2_IR_PD_DC photodiode and I found -0.6V. I tried to align it with the EQB2_GALVO_5 but was already centerred, Thus looking into the camera I realized that the beam was not in the usual position. I moved SQB2_M23 picomotors and I recovered the power -1.6V and the usual amplitude of the PDH signal. I performed 700 steps in horizontal and 350 in vertical
  4. I checked the SC PDH phase and was good but the sign wrt the green error signal was wrong thus I added pi rad to it. The new phase is 3.21 rad
  5. I checked the green lock UGF and was 70 Hz thus I increased the gain from 1 to 1.3

IR hand off

After the alignment I removed the IR normalization:

  • SC_PD_RF_NORM_SELECT was set to -1 i.e. forced to use the not normalized signal
  • I performed the TF between LFC_GR_RF_5MHz_NORM and LFC_SC_RF_11MHz_FIL and I found a factor 8.6. Thus I modified the filtering weight into the configuration file dividing it by 8.6. It is changed from 0.034759 to 0,0040417
  • I set the SC CND threshold to 5 for the rising and to 0.5 for the falling edge.

Then I started to perform some hand off. In total I performed 5 of them.

First Handoff 20:41 -> 20:47 UTC. I left the offload open (not checked why unlocked)

Second Handoff: 20:51 -> 20:57 UTC. Offload closed (lock trigger left on the green) (not checked why unlocked)

Third Handoff: 21:03 -> 21:13 UTC. Offload closed (lock trigger left on the green) It unlocked because I forgot to switch the error signal from SC to GR of the AOM loop but I swapped the sign. At a certain point the GR power dropped and it unlocked (figure 1). After this one I had to realigned in vertical the IR by 700 steps. Not understood why.

Fourth Handoff:  21:36 -> 21:39:36 UTC. Offload closed (lock trigger left on the IR) same mistake on the AOM loop but th trigger was on the IR, The unlock was caused by un unlock of the SC PLL

Fifth Handoff: 21:48 -> 22:23 UTC Here I solved all the mistake and I used the function for the handoff in the I solved a couple of bugs in it. The lock lasted more than half an hour but it unlocked again for the SC PLL unlock.

It seems tthat now the IR lock is stable. We need to understand why the SC PLL started to unlock time to time. 

Dufing the shift the Green power dropped from 7.7V to 6.5V. I found that the beam moved on the left in the FCEB_GR_CAM. I restored the usual power by moving the set point of the drift control. In particular all the set points now are 0 except FCEM_DRIFT_TY_SET = 12. Now the power is again 7.7V. Also the IR power dropped from 9.8V to 9 V. Restoring green power I restored also IR power.

I left the cavity locked on green with the IR in resonance and the AOM loop closed.

Concerning the I tested only the last version of the IRhandoff() and the GRrestore() functions. The GRhandoff() will be tested soon.


In Figure 3 in violet the spectrum of the error signals when we were locked on green and in blue when we were locked on IR. 

Images attached to this report
Comments to this report:
sorrentino, vardaro - 3:09 Monday 25 October 2021 (53631) Print this report

The reason why the SC AA was not working was because nonzero values were added to the error signals in the EQB1_HD_AA_dither.cfg file. Once corrected, the loop was engaged properly. We also updated the demodulation phase or SC_M5_X.

vardaro - 15:50 Monday 25 October 2021 (53637) Print this report

Additional details on the shift

  • after the third hand off the vertical control of SQB1 bench opened. That's why I had to realign the IR with SQB1 M21 V
  • I modified the offload logic: I multiplied the input signal with the lock flag and I modified the relay transition time from 100s to 1 s. In this way if the cavity unlocks the offload will open immediately without creating troubles.

Still to do:

  • save as SMS channel the weight of the LFC_Z_INPUT in order to have in the data the transitions from green to IR
  • Modify again the offload logic adding a second relay if we will discharge it again
  • Modify the drift control trigger. When we lock on the IR the drift control works with the IR but its trigger is still on the green

At the end of the shift (about 24UTC) I tried to change the proportional gain of the Slow loop of the SC PLL (from 40000 to 30000).  It seems that it worked. We will monitor this in the following days

Images attached to this comment
Search Help